The first part of the shift was dedicated to recover the problem of WEB - ADC7674 boards: Temperature too high; because of this all the suspension top stage boards received corrupted data at different times.
We decided to open all the suspensions loops and to investigated the problem of WEB ADC; I went at WE building and I found the fans of the crate 66 not working. The faulty crate has been changed with a spare one (see 51389).
After the replacement we could restore the communication from the WE; Paolo started the recovery of the susoensions.
At 9:48 UTC the cavities were locked; at 11:30 UTC we realized the SDB2 was controllable.
I contacted the SBE on call at 11:30 UTC that started the investigation from remote; the problem at SDB2 was finally fixed at 13:00 UTC and the ISC commissioning activity could restart.
this morning we experienced troubles with the name server; the problem seems related to the many instances of DataDisplay.
At 8:51 UTC the TB UPS1 switched off; I contacted Massimo and the UPS was switched on again. The reason of the switch off could be related to the high temperature in the room produced by a broken fan (see attached plot).
(12-04-2021 08:45 - 12-04-2021 09:00) On site
Description: TB UPS1 off
(12-04-2021 11:30 - 12-04-2021 13:00) From remote
Description: SDB2 not controllable
(12-04-2021 06:00 - 12-04-2021 09:48) From remote
Description: recovery of suspensions